# Michel F. C. Haddad

### v1

@@ 2 Identification in repeated cross-sections
(Conditional parallel trends):
$E[Y_{t}(d_{t}^{\prime})-Y_{t-1}(d_{t-1}^{\prime\prime})|D_{t}=d_{t},\mathbf{D_{t-1}=d_{t-1}^{\prime\prime}},X]\\ =E[Y_{t}(d_{t}^{\prime})-Y_{t-1}(d_{t-1}^{\prime\prime})|D_{t}=d_{t}^{\prime},\mathbf{D_{t-1}=d_{t-1}^{\prime\prime}},X]$
for all $d_{t}>d_{t}^{\prime}\geq 0$ in the support of $D_{t}$, $\mathbf{d_{t-1}^{\prime\prime}}\geq\mathbf{0}$ in the support of $\mathbf{D_{t-1}}$, and $t>0$ in the support of $T$.
Assumption 5 states that for two groups that receive a higher treatment dose $d_{t}$ (treatment group) and a lower dose $d_{t}^{\prime}$ (control group), respectively, in period $t$, and both receive treatment dose $d_{t-1}^{\prime\prime}$ in period $t-1$, the trend in mean potential outcomes when receiving dose $d_{t-1}^{\prime\prime}$ in period $t-1$ and dose $d_{t}^{\prime}$ in period $t$ is equal, conditional on $X$. For the special case where $d_{t}^{\prime}=d_{t-1}^{\prime\prime}$ as e.g., considered in de Chaisemartin et al. (2024), this implies that without changing the treatment between periods $t-1$ and $t$ for any group, the treatment and control groups would experience the same trend in their mean outcomes.
Assumption 5 is stronger than Assumption 1 due to the fact that it imposes parallel trends w.r.t. non-zero treatment doses, which generally requires further assumptions, such as homogeneity in average treatment effects across groups - see e.g., Fricke (2017). To illustrate this, assume that $\mathbf{D_{t-1}}$ is a single period (instead of a sequence) $D_{t-1}$, and consider setting (1) $d_{t}^{\prime}=d_{t-1}^{\prime\prime}=0$ (such that Assumption 5 is closely related to Assumption 1) and (2) $d_{t}^{\prime}>d_{t-1}^{\prime\prime}=0$. Taking differences in the conditional means in Assumption 5 under (2) and (1) demonstrates the implied effect homogeneity across groups with $D_{t}=d$ and $D_{t}=d^{\prime}$:
$E[Y_{t}(d_{t}^{\prime})-Y_{t-1}(0)|D_{t}=d_{t},D_{t-1}=0,X]-E[Y_{t}(0)-Y_{t-1}(0)|D_{t}=d_{t},D_{t-1}=0,X]$ (6)
$=$ $E[Y_{t}(d_{t}^{\prime})-Y_{t-1}(0)|D_{t}=d_{t}^{\prime},D_{t-1}=0,X]-E[Y_{t}(0)-Y_{t-1}(0)|D_{t}=d_{t}^{\prime},D_{t-1}=0,X]$
$\leftrightarrow$ $E[Y_{t}(d^{\prime})-Y_{t}(0)|D_{t}=d_{t},D_{t-1}=0,X]=E[Y_{t}(d^{\prime})-Y_{t}(0)|D_{t}=d_{t}^{\prime},D_{t-1}=0,X].$
@@
$f(D_{t}=d_{t},T=t|\mathbf{D_{t-1}},X,(D_{t},T)\in\{(d_{t}^{*},t^{\prime}),(d_{t},t)\})<1$
for all $d_{t}>0$ in the support of $D_{t}$, $t>0$ in the support of $T$, and $(d_{t}^{*},t^{\prime})\in\{(d_{t},t-1),(d_{t}^{\prime},t),(d_{t}^{\prime},t-1)\}$.
@@
$\Delta_{d_{t},d_{t}^{\prime},t}=E[Y_{t}(d_{t})-Y_{t}(d_{t}^{\prime})|D_{t}=d_{t}].$ (7)
It is worth noticing that under Assumptions 4 to 7, the conditional ATET of treatment dose $d_{t}$ versus $d^{\prime}_{t}$ in period $t$ given covariates $X$ and the previous treatment history $\mathbf{D_{t-1}}$ is analogously satisfied in expression (2), under the former Assumptions 1 to 4:
$E[Y_{t}(d_{t})-Y_{t}(d^{\prime}_{t})|D_{t}=d_{t},\mathbf{D_{t-1}},X]$
$=$ $E[Y_{t}|D_{t}=d_{t},\mathbf{D_{t-1}},X]-E[Y_{t-1}|D_{t}=d_{t},\mathbf{D_{t-1}},X]$
-- $\{E[Y_{t}|D_{t}=d^{\prime}_{t},\mathbf{D_{t-1}},X]-E[Y_{t-1}|D_{t}=d^{\prime}_{t},\mathbf{D_{t-1}},X]\}.$
Averaging the conditional ATET over the distribution of $X$ and $\mathbf{D_{t-1}}$ among the treated in any post-treatment period yields the ATET in that period:
@@
where $\mu_{d_{t}}(t^{\prime},\mathbf{d_{t-1}},x)=E[Y_{t^{\prime}}|D_{t}=d_{t},\mathbf{D_{t-1}}=\mathbf{d_{t-1}},X=x]$ is the conditional mean outcome in period $t^{\prime}$ given the treatment received in period $t$, the previous treatment history up to period $t-1$, and the covariates. We note that for the case that $d_{t}^{\prime}=d_{t-1}^{\prime\prime}$, the ATET corresponds to a causal effect of switching from the previous treatment dose $d_{t}^{\prime}$ to a higher treatment dose $d_{t}$ among switchers. As a comparison, de Chaisemartin et al. (2024) consider average effects among switchers across different treatment margins, rather than exclusively switching to dose $d_{t}$.
@@
As a final twist to our DiD framework, we might also be interested in the ATET of a treatment dose on lagged outcomes in later periods, rather than only evaluating the the impact of a treatment dose on an outcome in the exact same period. Hence, let us denote by $s\geq 1$ the time lag between the treatment and the outcome period, which permits defining the ATET of a treatment dose in period $t-s$ on an outcome in period $t$:
$\Delta_{d_{t-s},d^{\prime}_{t-s},t}=E[Y_{t}(d_{t-s})-Y_{t}(d^{\prime}_{t-s})|D_{t-s}=d_{t-s}],$ (11)
with time lag $s\geq 1$. We note that the ATET $\Delta_{d_{t-s},d^{\prime}_{t-s},t}$ comprises both the direct effect effect of the treatment dose in period $t-s$ on the outcome in $t$, as well as the indirect (or mediated) effect of the treatment dose in period $t-s$ on the treatments in later periods, that in turn also affect the outcome in period $t-s$. This is distinct from the scenario considered in De Chaisemartin and d’Haultfoeuille (2024), where the counterfactual is defined in terms of a fixed treatment dose $d^{\prime}$ over all post treatment periods $t-s$ to $t$, such that one could, for instance, consider the ATET $E[Y_{t}(d_{t-s})-Y_{t}(d^{\prime}_{t-s},...,d^{\prime}_{t})|D_{t-s}=d_{t-s}]$, in which any indirect effects of the treatment dose in period $t-s$ are absent.
$\Delta_{d_{t-s},d^{\prime}_{t-s},t}$ is identified when replacing Assumption 5 by the following parallel trend assumption:
@@
$E[Y_{t}(d_{t-s}^{\prime})-Y_{t-s-1}(d_{t-s-1}^{\prime\prime})|D_{t-s}=d_{t-s},\mathbf{D_{t-s-1}=d_{t-s-1}^{\prime\prime}},X]\\ =E[Y_{t}(d_{t-s}^{\prime})-Y_{t-s-1}(d_{t-s-1}^{\prime\prime})|D_{t-s}=d_{t-s}^{\prime},\mathbf{D_{t-s-1}=d_{t-s-1}^{\prime\prime}},X]$
for all $d_{t-s}>d_{t-s}^{\prime}\geq 0$ in the support of $D_{t-s}$, $\mathbf{d_{t-s-1}^{\prime\prime}}\geq\mathbf{0}$ in the support of $\mathbf{D_{t-s-1}}$, $t>1$ in the support of $T$, and $1\leq s<t$.
@@
Assumption 8 is stronger than Assumption 5, because it requires that changes in the treatment dose after period $t-s$ up to period $t$ do not evolve endogenously. This means that unobserved confounders jointly affecting the treatment doses in period $t-s$ and subsequent $s$ periods are ruled out. Otherwise, the parallel trend assumption would fail, as the outcome is also a function of the treatments after period $t-s$, such that the presence of unobserved confounders jointly affecting the treatment doses in period $t-s$ and subsequent periods would ultimately imply confounding of treatment $D_{t-s}$ and outcome $Y_{t}$. In contrast, Assumption 7 does not put restrictions on confounders of the treatment dose over time, with the caveat that the ATET may only be assessed on outcomes in the same period as the treatment dose to be evaluated, but not on outcomes in later periods.
@@
To identify the ATET under Assumptions 4, 6, 7, and 8 using an DR expression, we just need to appropriately modify equation (2) to include time lags when defining the pre- and and post-treatment periods of interest:
$\Delta_{d_{t-s},d^{\prime}_{t-s},t}=\lim_{h\rightarrow 0}E\left[\left\{\frac{\omega(D_{t-s};d_{t-s},h)\cdot I\{T=t\}}{\Pi_{d_{t-s},t}}\right.\right.$
×$\times[Y_{T}-\mu_{d_{t-s}}(t-s-1,\mathbf{D_{t-s-1}},X)-\mu_{d^{\prime}_{t-s}}(t,\mathbf{D_{t-s-1}},X)+\mu_{d^{\prime}_{t-s}}(t-s-1,\mathbf{D_{t-s-1}},X)]$
@@
Appendix A.1 demonstrates that DR expressions (2), (2), and (2) satisfy Neyman orthogonality. Consequently, those are attractive for constructing DML estimators under certain regularity conditions outlined in Chernozhukov et al. (2018).

### v2

@@ 3 Time-varying treatment doses in repeated cross-sections
(Conditional parallel trends with non-zero treatment doses):
$E[Y_{t}(d_{t}^{\prime})-Y_{t-1}(d_{t-1}^{\prime\prime})|D_{t}=d_{t},T=t,\mathbf{D}_{t-1}=d_{t-1}^{\prime\prime},X]\\ =E[Y_{t}(d_{t}^{\prime})-Y_{t-1}(d_{t-1}^{\prime\prime})|D_{t}=d_{t}^{\prime},T=t,\mathbf{D}_{t-1}=d_{t-1}^{\prime\prime},X]$
for $d_{t}>d_{t}^{\prime}\geq 0$ in the support of $D_{t}$, $\mathbf{d}_{t-1}^{\prime\prime}$ in the support of $\mathbf{D}_{t-1}$, and $t>0$ in the support of $T$.
Assumption 5 states that for two groups that receive a higher treatment dose $d_{t}$ (treatment group) and a lower treatment dose $d_{t}^{\prime}$ (control group), respectively, in period $t$, and both receive treatment dose $d_{t-1}^{\prime\prime}$ in period t1t-1, the trend in mean potential outcomes when moving from dose $d_{t-1}^{\prime\prime}$ in period t1t-1 to dose $d_{t}^{\prime}$ in period $t$ is the same across groups, conditional on $X$. For the special case where $d_{t}^{\prime}=d_{t-1}^{\prime\prime}$, see for example de Chaisemartin et al. (2024) and de Chaisemartin and D’Haultfœuille (2025), this implies that if no group changed the treatment dose between periods t1t-1 and $t$, then the treatment and control groups would experience the same trend in their mean outcomes. Moreover, we note that Assumption 5 is stronger than Assumption 1 because it imposes parallel trends with respect to non-zero treatment doses.
[…]
@@
[…]
@@
$\Delta_{d_{t},d_{t}^{\prime},t}=E[Y_{t}(d_{t})-Y_{t}(d_{t}^{\prime})|D_{t}=d_{t},T=t].$ (6)
It is worth noticing that under Assumptions 4 to 7, the conditional ATET of treatment dose $d_{t}$ versus $d^{\prime}_{t}$ in period $t$ given covariates $X$ and the previous treatment history $\mathbf{D}_{t-1}$ is identified analogously as expression (2):
$E[Y_{t}(d_{t})-Y_{t}(d^{\prime}_{t})|D_{t}=d_{t},T=t,\mathbf{D}_{t-1},X]$
$=$ $E[Y_{t}(d_{t})|D_{t}=d_{t},T=t,\mathbf{D}_{t-1},X]-E[Y_{t-1}(d_{t-1})|D_{t}=d_{t},T=t-1,\mathbf{D}_{t-1},X]$
- $\{E[Y_{t}(d_{t}^{\prime})|D_{t}=d^{\prime}_{t},T=t,\mathbf{D}_{t-1},X]-E[Y_{t-1}(d_{t-1})|D_{t}=d_{t},T=t-1,\mathbf{D}_{t-1},X]\}.$
Integrating the conditional ATET over the distribution of $X$ and $\mathbf{D}_{t-1}$ yields the ATET at a treatment dose $D_{t}=d_{t}$ in that period:
@@
where $\mu_{d_{t}}(t^{\prime},\mathbf{D}_{t-1},x)=E[Y_{T}|D_{t}=d_{t},T=t^{\prime},\mathbf{D}_{t-1}=\mathbf{D}_{t-1},X=x]$ is the conditional mean outcome in period $t^{\prime}$ given the treatment received in period $t$, the previous treatment history up to period t1t-1, and the covariates. We note that for the case that $d_{t}^{\prime}=d_{t-1}^{\prime\prime}$, the ATET corresponds to a causal effect of switching from the previous treatment dose $d_{t}^{\prime}$ to a higher treatment dose $d_{t}$ among switchers. As a comparison, de Chaisemartin et al. (2024) consider average effects among switchers across different treatment margins, rather than exclusively switching to dose $d_{t}$.
@@
As a further modification in our DiD framework, we might also be interested in the ATET of a lagged treatment dose on outcomes in later periods, rather than only evaluating the the instantaneous impact of a treatment dose on an outcome in the same period. Hence, let us denote by $s\geq 1$ the time lag between the treatment and the outcome period, which permits defining the ATET of a treatment dose in period tst-s on an outcome in period $t$:
$\Delta_{d_{t-s},d^{\prime}_{t-s},t}=E[Y_{t}(d_{t-s})-Y_{t}(d^{\prime}_{t-s})|D_{t-s}=d_{t-s},T=t],$ (10)
with time lag $s\geq 1$. We note that the ATET $\Delta_{d_{t-s},d^{\prime}_{t-s},t}$ comprises both the direct effect effect of the treatment dose in period tst-s on the outcome in $t$, as well as the indirect (or mediated) effect of the treatment dose in period tst-s on the treatments in later periods, that in turn also affect the outcome in period tst-s. $\Delta_{d_{t-s},d^{\prime}_{t-s},t}$ is identified when replacing Assumption 5 by the following parallel trend assumption:
@@
$E[Y_{t}(d_{t-s}^{\prime})-Y_{t-s-1}(d_{t-s-1}^{\prime\prime})|D_{t-s}=d_{t-s},T=t,\mathbf{D}_{t-s-1}=\mathbf{d}_{t-s-1}^{\prime\prime},X]\\ =E[Y_{t}(d_{t-s}^{\prime})-Y_{t-s-1}(d_{t-s-1}^{\prime\prime})|D_{t-s}=d_{t-s}^{\prime},T=t,\mathbf{D}_{t-s-1}=\mathbf{d}_{t-s-1}^{\prime\prime},X]$
for $d_{t-s}>d_{t-s}^{\prime}\geq 0$ in the support of $D_{t-s}$, $\mathbf{d}_{t-s-1}^{\prime\prime}\geq\mathbf{0}$ in the support of $\mathbf{D}_{t-s-1}$, $t>1$ in the support of $T$, and $1\leq s<t$.
@@
Assumption 8 is stronger than Assumption 5, because it requires that changes in the treatment dose after period tst-s up to period $t$ do not evolve endogenously. This means that unobserved confounders jointly affecting the treatment doses in period tst-s and subsequent $s$ periods are ruled out. Otherwise, the parallel trend assumption would generally fail, as the outcome (trend) is typically also a function of the treatments after period tst-s. Therefore, the presence of unobserved confounders jointly affecting the treatment doses in period tst-s and subsequent periods would ultimately imply confounding of treatment $D_{t-s}$ and outcome $Y_{t}$. In contrast, Assumption 5 does not put restrictions on confounders of the treatment dose over time, with the caveat that the ATET may only be assessed on outcomes in the same period as the treatment dose to be evaluated, but not on outcomes in later periods.
To better see the implications of Assumption 8 in terms of ruling out confounders that jointly affect treatments in different periods, consider the following example. Suppose the outcome in time period $T$ is a (possibly unknown) function, denoted by $\mathcal{F}_{Y}$, of the treatments in periods 1 and 2, $D_{1}$ and $D_{2}$, covariates $X$, and time $T$ (pre-treatment histories are omitted for simplicity), together with an additively separable, time-constant function $\mathcal{F}_{U}$ of time-invariant unobservables:
$Y_{T}=\mathcal{F}_{Y}(D_{1},D_{2},X,T)+\mathcal{F}_{U}(U).$
Further, assume that the first treatment is a function of $X$ and $U$,
$D_{1}=\mathcal{F}_{D_{1}}(X,U),$
and that the second treatment depends on the first treatment and $X$,
$D_{2}=\mathcal{F}_{D_{2}}(D_{1},X).$
We may add random, time-varying, and additively separable error terms to the models for $Y_{T}$, $D_{1}$, and $D_{2}$, but these are omitted here for simplicity.
Assume we are interested in the ATET of the treatment in period $T=1$ on the outcome in period $T=2$. Differencing average potential outcomes under treatment value $d_{1}^{\prime}$ between period $T=2$ and the pre-treatment period $T=0$, conditional on $X=x$, eliminates the time-constant component $\mathcal{F}_{U}(U)$ and yields
$E[Y_{2}(d^{\prime}_{1})-Y_{0}(d^{\prime}_{1})\mid X=x]=E[\mathcal{F}_{Y}(d^{\prime}_{1},D_{2}(d^{\prime}),x,2)-\mathcal{F}_{Y}(d^{\prime},D_{2}(d^{\prime}),x,0)],$ (11)
where $D_{2}(d_{1}^{\prime})$ denotes the potential second-period treatment under a first-period dose $d_{1}^{\prime}$.
Under the maintained model for $D_{2}$ (i.e., $D_{2}=\mathcal{F}_{D_{2}}(D_{1},X)$), the potential second-period treatment is $D_{2}(d_{1}^{\prime})=\mathcal{F}_{D_{2}}(d_{1}^{\prime},X)$ and hence does not depend on $U$. Consequently, $U$, while affecting $D_{1}$, does not affect the quantity on the right-hand side of (11); therefore, Assumption 8 is satisfied in this setup.
By contrast, if the second treatment also depends on $U$, say
$D_{2}=\mathcal{F}_{D_{2}}(D_{1},X,U),$
then the potential second-period treatment becomes $D_{2}(d_{1}^{\prime})=\mathcal{F}_{D_{2}}(d_{1}^{\prime},X,U)$, and $U$ enters the right-hand side of the differenced expression. In that case
$E[Y_{2}(d^{\prime}_{1})-Y_{0}(d^{\prime}_{1})\mid X=x]=E[\mathcal{F}_{Y}(d_{1}^{\prime},\mathcal{F}_{D_{2}}(d_{1}^{\prime},x,U),x,2)-\mathcal{F}_{Y}(d^{\prime},\mathcal{F}_{D_{2}}(d_{1}^{\prime},x,U),x,0)],$ (12)
so $U$ affects the mean potential outcome difference via the second treatment and also through the first treatment (since $D_{1}$ depends on $U$). Hence $U$ acts as a confounder for the effect of the first-period treatment on the differenced outcome, and Assumption 8 is violated. In summary, Assumption 8 rules out precisely such scenarios in which unobservables affect both the early-period treatment $D_{t-s}$ and subsequent treatments that also affect the outcome.
@@
To identify the ATET under Assumptions 4, 6, and 7 (when considering tst-s rather than $t$ as treatment period), as well as Assumption 8 using a DR expression, we need to appropriately modify equation (3) to include time lags when defining the pre- and post-treatment periods of interest:
[…]
$\times[Y_{T}-\mu_{d_{t-s}}(t-s-1,\mathbf{D}_{t-s-1},X)-\mu_{d^{\prime}_{t-s}}(t,\mathbf{D}_{t-s-1},X)+\mu_{d^{\prime}_{t-s}}(t-s-1,\mathbf{D}_{t-s-1},X)]$
@@
Appendix A.1 demonstrates that DR expressions (2), (3), and (3) satisfy Neyman orthogonality as the kernel bandwidth goes to zero. Therefore, these expressions can be employed to construct DML estimators under certain regularity conditions outlined in Chernozhukov et al. (2018).
We note that the identification of $\Delta_{d_{t-s},d^{\prime}_{t-s},t}$ under parallel trends as imposed in Assumption 8 is distinct from the scenario considered in De Chaisemartin and d’Haultfoeuille (2024), where the counterfactual is defined in terms of a fixed treatment dose $d^{\prime}$ over all post-treatment periods tst-s to $t$, that is, conditional on being an non-switcher in terms of treatment doses. This scenario allows considering the following ATET:
$\Delta_{d_{t-s},\mathbf{d^{\prime}}_{t-s:t},t}$ $=E[Y_{t}(d_{t-s})-Y_{t}(d^{\prime}_{t-s},...,d^{\prime}_{t})\mid D_{t-s}=d_{t-s},T=t]$ (14)
$=E[Y_{t}(d_{t-s})-Y_{t}(\mathbf{d^{\prime}}_{t-s:t})\mid D_{t-s}=d_{t-s},T=t],$
where $Y_{t}(\mathbf{d^{\prime}}_{t-s:t})$ denotes the lagged post-treatment counterfactual outcome under the sequence of fixed treatment doses $\mathbf{d^{\prime}}_{t-s:t}=(d^{\prime}_{t-s},\ldots,d^{\prime}_{t})$. Identification of the ATET in (14) requires yet another parallel trend assumption, different from Assumption 8 and more closely related to the one considered in De Chaisemartin and d’Haultfoeuille (2024):
Assumption 9
(Conditional parallel trends under a fixed treatment sequence):
$E\left[\,Y_{t}(d^{\prime}_{t-s},\ldots,d^{\prime}_{t})-Y_{t-s-1}(d^{\prime\prime}_{t-s-1})\mid D_{t-s}=d_{t-s},T=t,\mathbf{D}_{t-s-1}=\mathbf{d}^{\prime\prime}_{t-s-1},X\right]\\ =E\left[\,Y_{t}(d^{\prime}_{t-s},\ldots,d^{\prime}_{t})-Y_{t-s-1}(d^{\prime\prime}_{t-s-1})\mid D_{t-s}=d^{\prime}_{t-s},T=t,\mathbf{D}_{t-s-1}=\mathbf{d}^{\prime\prime}_{t-s-1},X\right]$
for $d_{t-s}>(d_{t-s}^{\prime},...,d^{\prime}_{t})\geq 0$ in the support of $(D_{t-s},...,D_{t})$, $\mathbf{d}_{t-s-1}^{\prime\prime}\geq\mathbf{0}$ in the support of $\mathbf{D}_{t-s-1}$, $t>1$ in the support of $T$, and $1\leq s<t$.
Assumption 9 requires that, conditional on covariates $X$ and the pre-period treatment history $\mathbf{D}_{t-s-1}$, the mean trend in potential outcomes under the sequence of fixed treatment doses $\mathbf{d}^{\prime}_{t-s:t}$ is the same for units whose realized treatment at time tst-s equals $d_{t-s}$ and for those with $d^{\prime}_{t-s}$. In contrast to Assumption 8, Assumption 9 allows for unobserved factors that jointly affect treatments in different time periods (provided that these factors do not also affect the differences in mean potential outcomes between the pre- and post-treatment periods). In the example given in equation (15), which entails the violation of Assumption 8, Assumption 9 is satisfied because, conditional on both the first- and second-period treatments, time-invariant unobserved confounders are differenced out from the mean potential outcome equations:
$E[Y_{2}(d^{\prime}_{1},d^{\prime}_{2})-Y_{0}(d^{\prime}_{1},d^{\prime}_{2})\mid X=x]=E[\mathcal{F}_{1}(d^{\prime}_{1},d^{\prime}_{2},x,2)-\mathcal{F}_{1}(d^{\prime}_{1},d^{\prime}_{2},x,0)].$ (15)
In addition to a modification of the parallel trend assumption, the nonparametric identification of the ATET in (14) requires strengthening the previous common support condition to hold with respect to the conditional density for the sequence:
Assumption 10
(Common support under a fixed treatment sequence):
[…]
This common support condition becomes more stringent as the number of post-treatment periods increases.
Under Assumptions 4, 6, 9, and 10, the ATET $\Delta_{d_{t-s},\mathbf{d^{\prime}}_{t-s:t},t}$ is identified by appropriately modifying DR expression (3). To account for a fixed treatment sequence $\mathbf{d}^{\prime}_{t-s:t}$, we define the product kernel
$\omega(\mathbf{D}_{t-s:t};\mathbf{d}^{\prime}_{t-s:t},\mathbf{h})=\prod_{r=0}^{s}\frac{1}{h_{r}}\,\mathcal{K}\!\left(\frac{D_{t-s+r}-d^{\prime}_{t-s+r}}{h_{r}}\right),$ (16)
where $\mathbf{h}=(h_{0},\ldots,h_{s})$ is a vector of bandwidths and $\mathcal{K}$ is a univariate kernel. Furthermore, we denote by $\mu_{\mathbf{d}^{\prime}_{t-s:t}}(t,\mathbf{D}_{t-s-1},X)=E[Y_{t}\mid\mathbf{D}^{\prime}_{t-s:t}=\mathbf{d}^{\prime}_{t-s:t},\mathbf{D}_{t-s-1},X]$ the conditional mean outcome under the treatment sequence, and by $\epsilon_{\mathbf{d}^{\prime}_{t-s:t},t,\mathbf{D}_{t-s-1},X}=Y_{t}-\mu_{\mathbf{d}^{\prime}_{t-s:t}}(t,\mathbf{D}_{t-s-1},X)$ the corresponding residual. Finally, the conditional density of the treatment sequence is given by $\rho_{\mathbf{d}^{\prime}_{t-s:t},t}(\mathbf{D}_{t-s-1},X)=f(\mathbf{D}^{\prime}_{t-s:t}=\mathbf{d}^{\prime}_{t-s:t}\mid\mathbf{D}_{t-s-1},X,T=t)$. Then, the following DR expression identifies the ATET defined in (14):
[…]
$\times[Y_{T}-\mu_{d_{t-s}}(t-s-1,\mathbf{D}_{t-s-1},X)-\mu_{\mathbf{d}^{\prime}_{t-s:t}}(t,\mathbf{D}_{t-s-1},X)+\mu_{\mathbf{d}^{\prime}_{t-s:t}}(t-s-1,\mathbf{D}_{t-s-1},X)]$
$-\left(\frac{\omega(\mathbf{D}_{t-s:t};\mathbf{d}^{\prime}_{t-s:t},\mathbf{h})\cdot I\{T=t\}\cdot\rho_{d_{t-s},t}(\mathbf{D}_{t-s-1},X)\cdot\epsilon_{\mathbf{d}^{\prime}_{t-s:t},t,\mathbf{D}_{t-s-1},X}}{\rho_{\mathbf{d}^{\prime}_{t-s:t},t}(\mathbf{D}_{t-s-1},X)\cdot\Pi_{d_{t-s},t}}\right$. (17)
$-\left.\left.\left.\frac{\omega(\mathbf{D}_{t-s:t};\mathbf{d}^{\prime}_{t-s:t},\mathbf{h})\cdot I\{T=t-s-1\}\cdot\rho_{d_{t-s},t}(\mathbf{D}_{t-s-1},X)\cdot\epsilon_{\mathbf{d}^{\prime}_{t-s:t},t-s-1,\mathbf{D}_{t-s-1},X}}{\rho_{\mathbf{d}^{\prime}_{t-s:t},t-s-1}(\mathbf{D}_{t-s-1},X)\cdot\Pi_{d_{t-s},t}}\right)\right\}\right].$
We note that in empircal applications, the use of product kernels in $(s+1)$ dimensions (related to the number of post-treatment periods) introduces the usual curse of dimensionality, since the effective sample size scales with $\prod_{r=0}^{s}h_{r}$.